OpenStudy (anonymous):

bernoulli equation..it will take a min to type out everything...solve 2ty'-y=2ty^3cost

OpenStudy (anonymous):

\[dy/dt-y/2t=y ^{3}cost\] divide y^3 from both sides \[y ^{-3}(dy/dt-y/(2t))=cost\] n=3, P=1/2t for standard form dy/dt+P(t)y=Q(t)y^n w=y^(1-n)=y^-2 dw/dt=-2y^-3dy/dt dy/dt=-y^3/2*dw/dt substitue dy/dt \[y ^{-3}(-y ^{3}dw/2dt-y/(2t))=cost\] \[-dw/2dt-1/(y ^{2}2t)=cost\] anything wrong so far? now integrating factor method... \[\mu(t)=e ^{\int\limits_{}^{}1/(2t)dt}=t ^{1/2}\] \[t ^{1/2}(-dw/2dt-w/(2t))=t ^{1/2} cost\] the w replaced 1/y^2 and now the part i'm stuck at...to find w, i have to integrate t^(1/2)cost...can't do it

OpenStudy (amistre64):

multiply it thru by : y^-3 \[ 2t y^{-3} y' - y^{-2} = 2t cost\] divide off the 2t \[ y^{-3} y' - (2t)^{-1}y^{-2} = cost\] substitute z = y^-2 z^-1/2 = y; \(\frac{-1}{2}z^{-3/2}z'=y'\) z^3/2 = y^-3 \[ \frac{-1}{2}z^{3/2}*z^{-3/2} \ z' - (2t)^{-1}z = cost\] \[ \frac{-1}{2}z' - (2t)^{-1}z = cost\] multiply by -2 \[ z' + (t)^{-1}z = -2cost\] and solve for z

OpenStudy (amistre64):

ive always used a "z" to name my substitute with

OpenStudy (amistre64):

when you multiply thru by that y^-n you automatically get you sub value of y^1-n

OpenStudy (anonymous):

i see where i went wrong, getting rid of the 2 made the difference thanks!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

i have a similar problem :( y'-2y=y^(-1/2)cost w=y^(3/2) y'=2/3y^(-1/2)w' w'-3w=3cost. next is int factor method.. mu(t)=e^(3t) w=(integral[e^(3t)3costdt]+C)/e^(3t) integrate e^(3t)3cost i must have done something wrong again

OpenStudy (amistre64):

\[y^{1/2}(y'-2y = y^{-1/2}cost)\] \[y^{1/2}y'-2y^{3/2} = cost\] \begin{align} z&=y^{3/2}\\ z^{2/3}&=y\implies &z^{1/3}=y^{1/2}\\ \frac{2}{3}z^{-1/3}z'&=y' \end{align} \[[z^{1/3}]\ [\frac{2}{3}z^{-1/3}z' ]-2[z] = cost\] \[\frac{3}{2}(\frac{2}{3}z'-2z = cost)\] \[z'-3z = \frac{3}{2}cost\]

OpenStudy (amistre64):

\[e^{-3x}(z'-3z = \frac{3}{2}cost)\] \[z'e^{-3x}-3e^{-3x}z = \frac{3}{2}e^{-3x}cost\] \[-3e^{-3x}z = \frac{3}{2}\int e^{-3x}cost\ dt\]

OpenStudy (anonymous):

stupid mistakes! thanks again, but i can't integrate that! well wolfram can but i have no idea what it did

OpenStudy (amistre64):

its integratable by parts and it should come back to itself so that you can "undo" it